# Parameter-Efficient Fine-Tuning of Large Language Models for Unit Test Generation: An Empirical Study

### v1

@@ Abstract
The advent of large language models (LLMs) like GitHub Copilot has significantly enhanced programmers’ productivity, particularly in code generation. However, these models often struggle with real-world tasks without fine-tuning. As LLMs grow larger and more performant, fine-tuning for specialized tasks becomes increasingly expensive. Parameter-efficient fine-tuning (PEFT) methods, which fine-tune only a subset of model parameters, offer a promising solution by reducing the computational costs of tuning LLMs while maintaining their performance. Existing studies have explored using PEFT and LLMs for various code-related tasks and found that the effectiveness of PEFT techniques is task-dependent. The application of PEFT techniques in unit test generation remains underexplored. The state-of-the-art is limited to using LLMs with full fine-tuning to generate unit tests. This paper investigates both full fine-tuning and various PEFT methods, including LoRA, (IA)3, and prompt tuning, across different model architectures and sizes. We use well-established benchmark datasets to evaluate their effectiveness in unit test generation. Our findings show that PEFT methods can deliver performance comparable to full fine-tuning for unit test generation, making specialized fine-tuning more accessible and cost-effective. Notably, prompt tuning is the most effective in terms of cost and resource utilization, while LoRA approaches the effectiveness of full fine-tuning in several cases.
Index Terms:
large language models, parameter-efficient fine-tuning, test generation, unit tests
I Introduction
@@
Large language models (LLMs) have been developed for various tasks, including general language and code-related activities, significantly enhancing programmers’ performance. Training LLMs on extensive code bases like GitHub give them seemingly magical capabilities. However, despite their proficiency in multiple code-related tasks, LLMs still struggle with cost-effectiveness in their training. The traditional approach to handling these specialized tasks has been fine-tuning the model specifically for each task. Given that LLMs often have billions, if not trillions, of parameters, traditional fine-tuning, which adjusts all parameters, is extremely computationally expensive and resource-intensive. However, state-of-the-art approaches [1] still rely on energy-costly full fine-tuning.
Advancements in instruction-tuned LLMs have encouraged investigations into various prompting and tuning techniques, e.g., [2, 3, 4], that do not require fine-tuning. This challenge has led to the development of multiple techniques for fine-tuning larger models with limited resources, commonly referred to as parameter-efficient fine-tuning (PEFT).
@@
The application of various PEFT methods has been explored for code-related generation tasks, such as code completion, code summarization, and more [5, 6, 7, 8, 9, 10, 11, 12, 13]. Evaluations of the PEFT methods show that they could achieve performance comparable to full fine-tuning while significantly reducing the computational burden by strategically fine-tuning only a chosen subset of parameters during the code-related generation process. However, the effectiveness of PEFT techniques on different tasks varies.
@@
LLMs have been used to generate unit test cases and assertion statements [14, 12, 2, 3, 4, 1]. However, to the best of our knowledge, the area of unit test generation using PEFT remains largely unexplored. Therefore, in this work, we empirically evaluate the performance of various PEFT methods for unit test generation. We focus on answering the following research questions:
@@
To answer the research questions, we compared full fine-tuning with three popular PEFT methods: LoRA (Low-Rank Adaptation) [15], (IA)3 (Infused Adapter by Inhibiting and Amplifying Inner Activations) [16], and prompt tuning [17]. We conducted these comparisons using ten LLMs from three open-source, decoder-only LLM families. The LLMs have varying architectures and sizes, ranging from 350 million to 16 billion parameters. We measured and compared the unit test generation performance of full fine-tuning and PEFT methods using well-established benchmark datasets rooted in real-world unit test cases.
Our results show that LoRA is generally the most reliable method for improving the quality of the generated unit tests and is, in many cases, on par with or better than full fine-tuning. Conversely, (IA)3 appears to be the least effective, often resulting in minimal improvements while consistently retaining prior knowledge in the LLMs. Prompt tuning demonstrates significant variability, excelling with some models, particularly larger ones, while underperforming with others. This variability highlights the importance of considering model size and characteristics when choosing the appropriate tuning method. Prompt tuning can be considered ”low-hanging fruit,” while LoRA should be employed for greater stability. Our contributions are as follows:
  • We conducted the first empirical study to extensively evaluate training LLMs using various PEFT methods—LoRA, (IA)3, and prompt tuning—for unit test generation across a wide range of LLM models.
@@
The rest of the paper is organized as follows. We provide preliminaries in Section II. Section III introduces related work. We explain our experimental design in Section IV. Section V presents the experimental results. Section VI discusses our results and limitations. Section VII concludes the study and proposes future work.
II Preliminaries
II-A LLM training
The classic language modeling task involves predicting the next word in a sequence, given the preceding words. The model is trained in an autoregressive manner. A typical example is the GPT model [18]. Training an autoregressive model involves fitting a function $f:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D}$, where $N$ represents the sequence length of tokens, each token having a dimension of $D$. The model aims to estimate the conditional probability distribution, denoted $p$, which can be expressed as:
$p(x_{1:T}\mid c)=\prod_{t=1}^{T}p(x_{t}\mid x_{1:t-1},c)$ (1)
where $x_{t}\in\mathbb{R}^{D}$ represents the $t$’th discrete token observation, $c$ denotes the initial context or input sequence, and $T$ signifies the length of the sequence of predicted tokens [19, Chapter 22].
LLMs are normally pre-trained on a huge amount of data, giving them a broad understanding of language patterns and semantics. This pre-training phase allows LLMs to capture intricate linguistic structures and contextual nuances, empowering them with a robust foundation for various natural language processing tasks.
Fine-tuning is a method in which a pre-trained model can be adapted to a more specific task, such as auto-completing source code. The process involves adjusting the weights of the pre-trained model using task-specific data, thereby enhancing its performance on a target task. When all parameters are tuned, this is often referred to as “full fine-tuning” [20]. If we imagine the model as a web of knowledge, the full fine-tuning loosens all the connections and lets them adapt entirely to the new domain. This flexibility allows for potentially superior performance but comes at a cost:
  • Overfitting: The model can become overly reliant on the training data, potentially struggling with unseen examples [21].
  • Computational cost: Training all parameters requires significant computational resources [20].
  • Catastrophic forgetting: The model might forget some valuable knowledge it acquired during pre-training [22, 23].
II-B PEFT Methods
Parameter-efficient fine-tuning (PEFT) methods provide an alternative way of fine-tuning a pre-trained model by training so-called adapters [24]. Here, most (if not all) parameters of the pre-trained model are frozen, and only a small number of parameters are trained. These adapters are often orders of magnitude smaller than the full model, making sharing, storing, and loading especially convenient. Moreover, some PEFT methods even facilitate the utilization of mixed-task batches, enabling distinct processing for different examples within a batch [17].
PEFT methods can be classified into three main classes of methods: additive, selective, and reparametrization-based [25]. Additive methods can be further broken down into adapters  [^1] The term ”adapter” is also used as the name of the first proposed adapter module for NLP. and soft prompts.
II-B1 Additive
This approach focuses on augmenting the pre-trained model with additional parameters or layers. Only the newly added components are trained, keeping the original model weights frozen. Popular techniques within the additive category are (IA)3 [16], adapters [26], and prompt tuning [17].
@@ II-B1 Additive
Adapter, or “bottleneck adapter,” introduced by Houlsby et al. [26] was arguably the first PEFT method. The class of “adapters” introduces small, fully connected networks placed after specific sub-layers within the model. These help the model adapt to the new task without significantly altering the core representation learned during pre-training.
@@
These techniques focus on creating the so-called “soft prompts.” Prompt tuning, introduced by Lester et al. [17], fine-tuned a subset of the model’s input embeddings, allowing the model to adapt to new tasks without altering the model’s parameters. The prefix tuning expands the idea to all model layers [27].
II-B2 Selective
@@
When using selective approaches, only a specific subset of the pre-trained model’s parameters is fine-tuned. This approach requires carefully selecting the parameters to be updated, ensuring that they are most relevant to the new task. An example is BitFit by Ben Zaken et al. [28], which only updates the bias parameters.
II-B3 Reparametrization-based
@@
This category utilizes techniques that reformulate the original model’s parameters into a more efficient trainable representation. This makes it possible to work with high-dimensional matrices while simultaneously reducing the number of trainable parameters. LoRA [15] is a common reparametrization-based method.
III Related Work
@@
This section presents state-of-the-art studies generating unit tests using LLMs and PEFT approaches to enhance code-related tasks.
III-A Unit test generation using LLMs
Researchers have turned to neural techniques to automatically generate unit tests to align with real-world best practices in unit test generation. Tufano et al. [1] propose an automated unit test case generation tool called AthenaTest. They construct a real-world focal method and developer-written test case dataset named Methods2Test 222 https://github.com/microsoft/methods2test and train a sequence-to-sequence transformer to generate unit test cases.
@@ III-A Unit test generation using LLMs
Advancements in instruction-tuned LLMs have also encouraged several investigations of various prompting techniques, as these do not require any fine-tuning [2, 3, 4]. Yuan et al. [3] propose ChatTester, a ChatGPT-based unit test generation approach, which leverages ChatGPT itself to improve the quality of its generated tests. Schäfer et al. [4] conducted a large-scale empirical evaluation on the effectiveness of LLMs for automated unit test generation without additional training or manual effort. Similarly to ChatTester, they also use ChatGPT (GPT-3.5-Turbo) and propose a re-prompting approach named TestPilot that automatically generates unit tests for all API functions in an npm package. Siddiq et al. [2] use the HumanEval and SF110 datasets to benchmark how well Codex, GPT-3.5-Turbo, and StarCoder can automatically generate unit tests. They evaluate the models based on compilation rates, test correctness, test coverage, and test smells.
III-B PEFT methods on code-related tasks
@@
While prompting methods are gaining traction, they often require significant manual effort to design complex templates and rely on expensive enterprise-grade LLMs like ChatGPT. Consequently, most cutting-edge methods for automatic code generation still involve computationally expensive full fine-tuning. One alternative approach to full fine-tuning is proposed by Steenhoek et al. [29]. They use Reinforcement Learning from Static Quality Metrics (RLSQM). As reinforcement learning is a very expensive fine-tuning method, they also see the value of using PEFT methods like LoRA to facilitate the training.
@@
PEFT has emerged as a powerful technique for enhancing code generation with LLMs. Researchers are investigating how these models perform on code tasks using PEFT techniques as a more efficient alternative to full fine-tuning and PEFT techniques’ effectiveness for different code-related tasks. Results show that the effectiveness of PEFT techniques is task-dependent and not transferrable. For example, Ayupov and Chirkova [5] evaluated adapters and LoRA on various tasks, finding them effective for code understanding but less so for code generation compared to full fine-tuning. Wang et al. [7] reports that prompt tuning of CodeBERT and CodeT5 outperforms full fine-tuning for defect prediction, code summarization, and code translation tasks. Liu et al. [13] evaluated adapter, LoRA, Prefix tuning, and Mix-And-Match adapter (MAM) [30] for defect detection, code clone detection, and code summarization. They concluded that “PEFT may achieve comparable or higher performance than full fine-tuning in code understanding tasks, but they may exhibit slightly weaker performance in code generation tasks.” Similarly, Weyssow et al. [8] evaluate text-to-code generation using LoRA, (IA)3, prompt tuning, and prefix tuning. They compare performance across models of varying sizes (up to 7 billion parameters) against full fine-tuning of smaller models (less than 350 million parameters) and concluded that LoRA is the most effective one.
@@
To our knowledge, there is currently no systematic knowledge of how various PEFT methods behave in the domain of unit test generation. As existing studies show that PEFT techniques’ effectiveness is task-dependent, it is, therefore, necessary to empirically evaluate their effectiveness for unit test generation.
IV Experimental design
@@
The main objective of this study is to gain insight into the performance of fine-tuning LLMs using PEFT for unit test generation.
IV-A Research Questions
@@
To address RQ1 (”How well does PEFT perform on unit test generation?”), we evaluate several different PEFT methods on unit test generation. We also compare the PEFT methods with full fine-tuning. Our study covers a wide range of language models for code of various sizes, spanning 350 million to 16 billion parameters. We assess the quality of the generated test cases produced by different PEFT methods by scoring against real-world human-written reference tests and evaluating the test cases’ syntactic correctness. Further, we examine whether applying PEFT methods for unit test generation adversely affects the pre-trained models’ code generation performance, also known as catastrophic forgetting [22, 23]. This evaluation will help us understand if there are any trade-offs between applying fine-tuning and PEFT methods for unit test generation and maintaining performance on other code-related tasks.
@@
Regarding RQ2 (”What is the relation between resource utilization and performance of PEFT in unit test generation?”), we want to understand PEFT methods’ performance gains in relation to their resource utilization and requirements. Thus, we analyze the efficiency of PEFT methods in terms of the computational resources they use and their impact on the performance outcomes in unit test generation.
IV-B Fine-tuning methods selection

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) methods, which fine-tune only a subset of model parameters, offer a promising solution by reducing the computational costs of tuning large language models (LLMs) while maintaining their performance. Existing studies have explored using PEFT and LLMs for various code-related tasks and found that the effectiveness of PEFT techniques is task-dependent. The state-of-the-art is limited to using LLMs with full fine-tuning to generate unit tests. The application of PEFT techniques in unit test generation remains underexplored. This paper investigates both full fine-tuning and various PEFT methods, including LoRA, (IA)3, and prompt tuning, across thirteen models of different architectures and sizes. We use well-established benchmark datasets to evaluate their effectiveness in unit test generation and measure syntax correctness, CodeBLEU, pass@1, instruction coverage, branch coverage, and mutation score of the generated tests. Our findings show that LoRA can deliver performance comparable to full fine-tuning for unit test generation in several cases. If training costs are valued, prompt tuning is the most cost-effective approach, particularly for large models. However, the models tuned with full fine-tuning or PEFT may generate fewer executable test cases than the baseline model because they generate more tests calling nonexistent methods or having type mismatches. For the generated ones that are executable, the ones from the tuned models show better test coverage than those from the baseline model.
1 Introduction
@@
Training LLMs on extensive code bases like GitHub gives them seemingly magical capabilities. However, despite their proficiency in multiple code-related tasks, LLMs still struggle with the high cost of their training. The traditional approach to handling specialized tasks has been fine-tuning the model for each task. Given that LLMs often have billions, if not trillions, of parameters, traditional fine-tuning, which adjusts all parameters, is extremely computationally expensive and resource-intensive [tufano2021unit].
@@
Advancements in LLMs have encouraged investigations into various prompting and tuning techniques, e.g., [siddiq2024using, yuan2023manual, schäfer2023empirical], that do not require full fine-tuning. This challenge has led to the development of multiple techniques for fine-tuning larger models with limited resources, commonly referred to as parameter-efficient fine-tuning (PEFT). The application of various PEFT methods has been explored for code-related tasks, such as code completion, code summarization, and more [Ayupov2022ParameterEfficientFO, Goel2022OnTC, Wang2023PromptTI, weyssow2024exploring, Wang2023OneAF, Shi2023TowardsEF, Liu2023MFTCoderBC, yuan2023evaluating, Liu2023AnES]. Evaluations of the PEFT methods show that they could achieve performance comparable to full fine-tuning while significantly reducing the computational burden by strategically fine-tuning only a chosen subset of parameters for code-related tasks. However, the effectiveness of PEFT techniques on different tasks varies [Ding2023ParameterefficientFO].
@@
LLMs have been used to generate unit test cases and assertion statements [Mastropaolo2021StudyingTU, yuan2023evaluating, siddiq2024using, yuan2023manual, schäfer2023empirical, tufano2021unit]. However, to the best of our knowledge, the area of empirical comparisons of PEFT and full fine-tuning for unit test generation remain unexplored. Therefore, in this work, we empirically evaluate the performance of various PEFT methods for unit test generation. We focus on answering the following research questions:
@@
To answer the research questions, we compared full fine-tuning with three popular PEFT methods: LoRA (Low-Rank Adaptation) [Yu2023LowRankAO], (IA)3 (Infused Adapter by Inhibiting and Amplifying Inner Activations) [Liu2022FewShotPF], and prompt tuning [Lester2021ThePO]. We conducted the comparisons using thirteen LLMs from four open-source decoder-only LLM families. The LLMs have varying architectures and sizes, ranging from 350 million to 16 billion parameters. We measured and compared the unit test generation performance of full fine-tuning and PEFT methods using well-established benchmark datasets rooted in real-world unit test cases and metrics, including syntactic correctness, CodeBLEU, pass@1, instruction coverage, branch coverage, and mutation score.
Our results show that LoRA is generally the most reliable PEFT method to generate good-quality unit tests and is, in many cases, on par with or even better than full fine-tuning. Regarding cost-effectiveness, prompt tuning provides the best results, particularly for larger models. Prompt tuning can be considered “low-hanging fruit”, while LoRA should be employed for greater performance. The results also reveal that the generated unit tests from tuned models can contain more compilation errors, which need to be investigated in future studies. Our contributions are as follows:
  • We conducted the first empirical study to extensively evaluate tuning LLMs using various PEFT methods—LoRA, (IA)3, and prompt tuning—for unit test generation across a wide range of LLM models.
@@
The rest of the paper is organized as follows. Section 2 introduces the PEFT methods we explore. Section 3 introduces related work. We explain our experimental design in Section 4. Section 5 presents the experimental results. Section 6 discusses our results and limitations. Section 7 concludes the study and proposes future work.
2 PEFT Methods
Parameter-efficient fine-tuning (PEFT) methods provide an alternative way of fine-tuning a pre-trained model by training so-called adapters [Ding2023ParameterefficientFO]. Here, most (if not all) parameters of the pre-trained model are frozen, and only a small number of parameters are trained. These adapters are often orders of magnitude smaller than the full model, making sharing, storing, and loading especially convenient. Moreover, some PEFT methods even facilitate the utilization of mixed-task batches, enabling distinct processing for different examples within a batch [Lester2021ThePO]. PEFT methods can be classified into three main categories: additive, selective, and reparametrization-based [lialin2023scaling]. Additive methods can be further broken down into adapters  [^1] The term “adapter” is also used as the name of the first proposed adapter module for NLP. and soft prompts.
2.1 Additive
This approach focuses on augmenting the pre-trained model with additional parameters or layers. Only the newly added components are trained, keeping the original model weights frozen. Popular techniques within the additive category are (IA)3 [Liu2022FewShotPF], adapters [houlsby2019parameter], and prompt tuning [Lester2021ThePO].
@@ 2.1 Additive
Adapter, or “bottleneck adapter,” introduced by [houlsby2019parameter], was arguably the first PEFT method. The class of “adapters” introduces small, fully connected networks placed after specific sub-layers within the model. These help the model adapt to the new task without significantly altering the core representation learned during pre-training.
@@
These techniques focus on creating the so-called “soft prompts.” Prompt tuning, introduced by [Lester2021ThePO], fine-tuned a subset of the model’s input embeddings, allowing the model to adapt to new tasks without altering the model’s parameters. The prefix tuning expands the idea to all model layers [li-liang-2021-prefix].
2.2 Selective
@@
When using selective approaches, only a specific subset of the pre-trained model’s parameters is fine-tuned. This approach requires carefully selecting the parameters to be updated, ensuring that they are most relevant to the new task. An example is BitFit by [BenZaken2021BitFitSP], which only updates the bias parameters.
2.3 Reparametrization-based
@@
This category utilizes techniques that reformulate the original model’s parameters into a more efficient trainable representation. This makes it possible to work with high-dimensional matrices while simultaneously reducing the number of trainable parameters. LoRA [Yu2023LowRankAO] is a typical reparametrization-based method.
3 Related Work
@@
This section presents state-of-the-art studies generating unit tests using LLMs and studies applying PEFT approaches to enhance code-related tasks.
3.1 Unit test generation using LLMs
@@
[tufano2021unit] propose an automated unit test case generation tool called AthenaTest. They construct a real-world focal method and developer-written test case dataset named Methods2Test  [^2] https://github.com/microsoft/methods2test and train a sequence-to-sequence transformer to generate unit test cases. [yuan2023manual] propose ChatTester, a ChatGPT-based unit test generation approach, which leverages ChatGPT itself to improve the quality of its generated tests. [schäfer2023empirical] conducted a large-scale empirical evaluation on the effectiveness of LLMs for automated unit test generation without additional training. Similarly to ChatTester, they also use ChatGPT (GPT-3.5-Turbo) and propose a re-prompting approach named TestPilot. TestPilot automatically generates unit tests for all API functions in an npm package, achieving a median statement and branch coverage of 70.2% and 52.8% on 1,684 API functions, respectively. [siddiq2024using] use the HumanEval and SF110 datasets to benchmark how well Codex, GPT-3.5-Turbo, and StarCoder can automatically generate unit tests. They evaluate the models based on compilation rates, test correctness, test coverage, and test smells. [alagarsamy2024a3test] propose A3Test, a deep-learning based unit-test generation method that augments test generation with assertion knowledge and verifies naming and signatures. They evaluated the method on 5,278 focal methods from the Defects4J [just2014defects4j] benchmark. The methods by [yuan2023manual, siddiq2024using, schäfer2023empirical, alagarsamy2024a3test] primarily leverage large off-the-shelf LLMs, which can be costly and often lack the flexibility for custom fine-tuning to generate unit tests for specific domains and organizations. Additionally, such prompting-based methods often require significant manual effort to design complex templates and rely on expensive enterprise-grade LLMs like ChatGPT.
3.2 PEFT methods on code-related tasks
@@
Many cutting-edge methods for automatic code generation still involve computationally expensive full fine-tuning. One alternative approach to full fine-tuning is proposed by [steenhoek2023reinforcement]. They use Reinforcement Learning from Static Quality Metrics (RLSQM). As reinforcement learning is a very expensive fine-tuning method, they also see the value of using PEFT methods like LoRA to facilitate the training.
@@
PEFT has emerged as a powerful technique for enhancing code generation with LLMs. Researchers are investigating how to apply PEFT techniques to code-related tasks as a more efficient alternative to full fine-tuning. Results show that the effectiveness of PEFT techniques is task-dependent and not transferable. For example, [Ayupov2022ParameterEfficientFO] evaluated adapters and LoRA on various tasks, finding them effective for code understanding but less so for code generation compared to full fine-tuning. [Wang2023PromptTI] report that prompt tuning of CodeBERT and CodeT5 outperforms full fine-tuning for defect prediction, code summarization, and code translation tasks. [Liu2023AnES] evaluated adapter, LoRA, Prefix tuning, and Mix-And-Match adapter (MAM) [mam] for defect detection, code clone detection, and code summarization. They concluded that “PEFT may achieve comparable or higher performance than full fine-tuning in code understanding tasks, but they may exhibit slightly weaker performance in code generation tasks.” Similarly, [weyssow2024exploring] evaluated text-to-code generation using LoRA, (IA)3, prompt tuning, and prefix tuning. They compared performance across models of varying sizes (up to 7 billion parameters) against full fine-tuning of smaller models (less than 350 million parameters) and concluded that LoRA is the most effective one.
@@
As existing studies show that PEFT techniques’ effectiveness is task-dependent, it is, therefore, necessary to empirically evaluate their effectiveness for unit test generation. To our knowledge, there is currently no systematic knowledge of how various PEFT methods behave to tune LLMs for unit test generation. In this study, we focus on autoregressive base code models to minimize the influence of prompt templates on PEFT performance.
4 Experimental design
@@
The main objective of this study is to gain insight into the performance and cost-effectiveness of fine-tuning LLMs using PEFT for unit test generation.
4.1 Research Questions
@@
To answer RQ1 (“How well does PEFT perform on unit test generation?”), we aim to evaluate several different PEFT methods on unit test generation. We also want to compare the PEFT methods with full fine-tuning. Thus, our study covers a wide range of language models for code of various sizes, spanning 350 million to 16 billion parameters. We assess the quality of the generated test cases by different PEFT methods by scoring against real-world human-written reference tests and evaluating the test cases’ syntactic correctness, similarity to reference tests, passing rates, instruction and branch coverages, and mutation scores. Further, we examine whether applying PEFT methods for unit test generation adversely affects the pre-trained models’ code generation performance, also known as catastrophic forgetting [French1999CatastrophicFI, Luo2023AnES]. Organizations may use the same model to generate both code and unit tests within their integrated development environments (IDEs). This evaluation will help us understand if there are any trade-offs between applying fine-tuning and PEFT methods for unit test generation and maintaining performance on other code-related tasks.
@@
To answer RQ2 (“What is the relation between resource utilization and performance of PEFT in unit test generation?”), we analyze the efficiency of PEFT methods in terms of the computational resources they use and the performance they achieve in unit test generation.
4.2 Fine-tuning methods selection